OpenStudy (anonymous):

Find the slope of the tangent line equation of the curve x^2 lny+xy^3= 5 at the point (5,1) on the curve by using implicit differentiation

OpenStudy (anonymous):

\[\begin{align*} x^2\ln y+xy^3&=5\\\\ \frac{d}{dx}\left[x^2\ln y+xy^3\right]&=\frac{d}{dx}5\\\\ \frac{d}{dx}\left[x^2\ln y\right]+\frac{d}{dx}\left[xy^3\right]&=0\\\\ x^2\frac{d}{dx}\left[\ln y\right]+\ln y\frac{d}{dx}\left[x^2\right]+y^3\frac{d}{dx}\left[xy^3\right]+x\frac{d}{dx}\left[y^3\right]&=0\\\\ \cdots&=0 \end{align*}\] Any problems so far?

OpenStudy (freckles):

\[\frac{d}{dx}(xy^3)=x \frac{d}{dx}(y^3)+y^3 \frac{d}{dx}(x)\]